# Bridging Developer Needs and Feasible Features for AI Assistants in IDEs

### v2

@@ Abstract
Despite the increasing presence of AI assistants in Integrated Development Environments, it remains unclear what developers actually need from these tools and which features are likely to be implemented in practice. To investigate this gap, we conducted a two-phase study. First, we interviewed 35 professional developers from three user groups (Adopters, Churners, and Non-Users) to uncover unmet needs and expectations. Our analysis revealed five key areas: Technology Improvement, Interaction, and Alignment, as well as Simplifying Skill Building, and Programming Tasks. We then examined the feasibility of addressing selected needs through an internal prediction market involving 102 practitioners. The results demonstrate a strong alignment between the developers’ needs and the practitioners’ judgment for features focused on implementation and context awareness. However, features related to proactivity and maintenance remain both underestimated and technically unaddressed. Our findings reveal gaps in current AI support and provide practical directions for developing more effective and sustainable in-IDE AI systems.
Index Terms:
@@
Human-AI Collaboration, Artificial Intelligence, Generative AI, LLMs, Intelligent Assistants, Integrated Development Environment, Programming
I INTRODUCTION
@@
The increasing integration of Artificial Intelligence (AI) tools in software development is transforming the way developers interact with their Integrated Development Environments (IDEs) [1, 2, 3]. These AI-powered assistants facilitate tasks like code generation, debugging, and optimization with the goal of boosting productivity and reducing the cognitive load associated with repetitive programming tasks [4, 5, 6, 7]. As AI-based systems evolve rapidly, teams that build tools for tech creators face two key challenges: understanding how coding specialists actually use AI inside their IDEs and deciding which features are worth investing in.
@@
However, actionable insights into in-IDE human-AI experience (HAX) are still limited. While recent studies have explored developers’ general expectations for AI-assisted tools, few focus specifically on in-IDE interactions, particularly the successes and challenges developers face and their needs with these tools [7]. Even fewer studies consider differences across user groups: those who adopted AI into their workflow, those who abandoned it, and those who have never tried it. These distinctions are important for informing development and marketing strategies [8]. Moreover, these studies often focus on usage patterns or user perceptions without connecting them to concrete feature development decisions. As a result, we know little about how specific user needs translate into buildable and valuable AI features.
@@
To address existing knowledge gaps, we conducted a multi-method study in two phases: qualitative semi-structured interview [9] and prediction market [10]. First, we interviewed 35 tech creators across three usage groups (Adopters, Churners, and Non-users) to gain qualitative insights into their experiences, challenges, and unmet needs with current AI features. This phase reveals unmet needs and expectations across five areas: Technology Improvement, Interaction, Alignment, Simplifying Skill Building, and Programming Tasks. Yet, understanding user needs alone does not directly translate into feasible product development strategies. Prioritizing feature development requires additional insight into both technical feasibility and organizational readiness. To achieve this, we complemented our qualitative findings with a prediction market, an approach in which participants make predictions on future events, leveraging collective judgment to estimate outcomes. We asked 102 internal practitioners to predict the likelihood of implementing 12 user-desired features. This allows us to integrate user-driven insights with practical considerations from the very people who shape our product development. Our results reveal that internal practitioners correctly anticipate the implementation of features around AI in Implementation and Context-aware Technology. In contrast, features requiring proactive behavior and maintenance support remain undervalued and underdelivered. These findings highlight a misalignment between developer needs and current feasibility assessments, pointing to concrete opportunities for more helpful and forward-looking AI assistance in IDEs.
@@
With this work, we aim to provide industry practitioners and researchers with structured evidence-based insights to guide product decisions. Specifically, offer:
@@
  • A categorization of user needs for in-IDE HAX to outline the technological improvements needed to make AI tools more useful;
@@ II BACKGROUND
AI-powered programming assistants have become increasingly common, with tools such as GitHub Copilot, Cursor, and JetBrains AI integrated into developer workflows [4, 5, 6]. While their technical capabilities continue to evolve, research into how developers experience and evaluate such tools has also grown.
Recent studies have investigated what professional developers need and expect from AI-powered programming assistants [11, 12, 13, 14, 15, 16, 17]. Across contexts, developers consistently seek tools that reduce repetitive effort and provide relevant suggestions without introducing new debugging burdens [11, 16, 15]. Effective assistance is expected to align with the current task, understand code context, and support natural language interaction [11, 13, 18].
Transparency and control are also emphasized, including the ability to steer and correct outputs and control context [12, 11]. Several studies highlight that experienced users tend to expect more than raw generation [13, 12]. For example, tools are valued when they adapt to user intent and surface relevant, task-specific information [11, 13, 19].
At the same time, integration challenges persist. Developers often report friction when tools disrupt their workflow or fail to fit specific environments [14, 15]. Outside of initial implementation stages, non-functional requirements such as performance and scalability may be harder to maintain [14]. These insights point to the need for assistants that can flexibly support diverse roles, offer customization, and operate within existing development processes [12, 14, 11].
@@
Despite these insights, most studies examine developer interaction with AI in broad or external settings, without addressing the unique characteristics of IDE-based workflows. This distinction is important. According to large-scale industry reports, up to 97% of developers rely on IDEs for software tasks [20]. In-IDE AI assistants operate within a structured environment that offers richer contextual signals such as file scope, project layout, and cursor history. These tools have the potential to deliver more context-aware support while minimizing context switching. Yet, their design introduces new constraints, including the need to remain non-intrusive and responsive in real time. Therefore, in-IDE AI tools shape distinct user needs, expectations, and challenges that warrant focused investigation.
@@
Moreover, existing studies tend to treat developers as a single user group, overlooking differences across adoption stages [7]. However, industry practice suggests that design and feature prioritization must account for the heterogeneous needs of distinct groups: Adopters, Churners, and Non-Users [8, 21]. These groups not only engage (or not) with AI tools differently, but also hold divergent mental models, trust levels, and expectations, which may lead to different product requirements or marketing strategies. Recognizing these “personas” is necessary to design AI assistants that support retention, re-engagement, or onboarding.
@@
At the same time, even when user needs are clearly articulated, translating them into product decisions remains a challenge. Feature roadmaps must also reflect internal constraints such as technical feasibility, resource allocation, and infrastructure readiness. Traditional prioritization techniques, such as expert panels and surveys, often fall short by providing limited collective insights or biased individual opinions [22]. An alternative approach, prediction markets, has been shown to effectively aggregate collective expert judgment to forecast the likelihood of specific outcomes [10]. A prediction market is a structured method wherein participants buy and sell predictions about future events, effectively leveraging the “wisdom of the crowd” to estimate probabilities accurately [10, 23].
@@
Thus, this paper combines qualitative interviews and prediction markets, integrating rich user insights with realistic assessments from internal stakeholders. This dual perspective enables a grounded understanding of both user needs and internal feasibility, offering a practical contribution to the design and deployment of in-IDE AI assistants.
III UNDERSTANDING DEVELOPER NEEDS
In this section, we examine how developers experience AI inside their IDEs when coding. Our goal is to provide a structured understanding of what developers need from AI assistance, what challenges they face, and how these vary across adoption groups.
We address the following research questions:
  • RQ1. What are developers’ concrete needs when using AI tools within IDEs?
@@ III UNDERSTANDING DEVELOPER NEEDS
III-A METHOD
@@
The study was conducted in line with our institutional ethical standards, adhering to the values and guidelines outlined in the ICC/ESOMAR International Code [24].
III-A1 Recruitment
@@
We recruited participants from users of JetBrains, a software engineering company that also develops AI-based tools for software development, to ensure a consistent experience with development environments. Note that with that, participants could use any in-IDE AI assistant. Invitations were sent via email to consenting subscribers and promoted through the authors’ social media.
Respondents completed a qualification survey designed to assign them to one of three groups: Adopters, Churners, and Non-Users. We chose these three groups to capture various experiences with AI tools in IDEs [8]. These groups are referred to using the prefixes A, C, and N, respectively, in participant identifiers throughout the paper.
In the survey, we collected data on job roles, experience levels, IDEs used, AI tools tried for coding, and duration of AI usage. Out of 380 responses, we iteratively invited participants until data saturation was reached [25], which means that no new themes or insights emerged from the last interviews. As a result, we conducted 35 interviewees: 15 Adopters, 12 Churners, and 8 Non-Users.
Most of the participants were developers (32) with overlapping roles: 9 team leads, 8 architects, and 5 DevOps engineers. The sample included 28 seniors, 5 mid-level practitioners, and 2 juniors, with experience ranging from under 5 years (9) to over 11 years (16). Participants chose a reward between a USD 100 Amazon gift card or an equivalent-value pack of all JetBrains products.
III-A2 Data Collection
@@
The interviews followed a semi-structured script developed by the first two authors, reviewed by certified UX researchers at JetBrains, and piloted with four internal participants [9]. The script included open-ended questions on AI experience, utilized tools, challenges and successes with in-IDE HAX, unmet needs, desired features, and visions for AI in coding.
@@
To account for differing usage patterns, we tailored the script per group. For Adopters, the questions focused on their ongoing use of AI tools, exploring typical tasks, workflows, challenges, and successes. Churners reflected on their initial adoption and further disengagement, identifying reasons for discontinuation and missed opportunities. Non-Users were asked about their reasons for not adopting AI, any potential barriers and possible motivations that could encourage them to begin using these tools. This ensured group-specific depth while maintaining thematic consistency across interviews.
III-A3 Data Analysis
All interviews were transcribed using Condens [^1] Condens: a user research repository for analyzing qualitative data. https://condens.io/. We applied a thematic analysis [26, 27] to identify user-reported successes, challenges, unmet needs, and future expectations for in-IDE AI.
@@ III-A3 Data Analysis
The first two authors reviewed the first transcript together to agree on the analysis and the code-deriving process. They then went into an iterative inductive process, where they independently assigned Codes to quotes from the transcripts. Eventually, they compiled a list of 161 codes, met again to add or remove codes, and resolved any disagreements until full consensus was reached. The final versions of the codes were then assigned to the corresponding quotes. The codes were then grouped into Thematic groups representing key areas of need. These were organized into broader Topics that describe the categories of in-IDE HAX needs.
For example, the following quote:
@@
was coded as “Code Review Support”, assigned to the group “Code Optimization”, under the topic “Simplifying Programming Tasks”.
@@
All materials used for data gathering, together with a full demography table of participants and a summarized codebook, are included in Supplementary Materials [28]. Full transcripts of interviews cannot be shared due to the company’s confidentiality policy.
III-B FINDINGS
@@
Figure 1: The Categorization of Users’ Needs in HAX Within IDEs With Descriptions

### v3

# Developer Needs and Feasible Features for AI Assistants in IDEs

@@ Abstract.
Despite the increasing presence of AI assistants in Integrated Development Environments (IDEs), it remains unclear what different groups of developers actually need from these tools and which features are likely to be implemented in practice. To investigate this gap, we conducted a two-phase study. First, we interviewed 35 professional developers from three user groups (Adopters, Churners, and Non-Users) to uncover unmet needs and expectations. Our analysis revealed five key areas of need distinctly distributed across practitioners’ groups: Technology Improvement, Interaction, and Customization, as well as Simplifying Skill Building, and Programming Tasks. We then examined the feasibility of addressing selected needs through an internal prediction market involving 102 practitioners. The results demonstrate a strong alignment between the developers’ needs and the practitioners’ judgment for features focused on implementation and context awareness. However, features related to proactivity and maintenance remain both underestimated and technically unaddressed. Our findings reveal gaps in current AI support and provide practical directions for developing more effective and sustainable in-IDE AI systems.
@@
Human-AI Collaboration, Artificial Intelligence, Generative AI, LLMs, Intelligent Assistants, Integrated Development Environment
copyright: nonejournalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIP ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786583.3786883isbn: 979-8-4007-2426-8/2026/04ccs: Human-centered computing User studiesccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Integrated and visual development environments
1. INTRODUCTION
@@
The increasing integration of Artificial Intelligence (AI) tools in software development is transforming the way developers interact with their Integrated Development Environments (IDEs) (Marron, 2024; Izadi et al., 2024; Semenkin et al., 2024). These AI-powered assistants facilitate tasks like code generation, debugging, and optimization with the goal of boosting productivity and reducing the cognitive load associated with repetitive programming tasks (12; 16; Cursor AI (2024); A. Sergeyuk, I. Zakharov, E. Koshchenko, and M. Izadi (2025)). As AI-based systems evolve rapidly, teams that build tools for tech creators face two key challenges: understanding how coding specialists actually use AI inside their IDEs and deciding which features are worth investing in.
@@
However, actionable insights into in-IDE human-AI experience (HAX) are still limited. While recent studies have explored developers’ general expectations for AI-assisted tools, few focus specifically on in-IDE interactions, particularly the successes and challenges developers face and their needs with these tools (Sergeyuk et al., 2025). Even fewer studies consider differences across user groups: those who adopted AI into their workflow, those who abandoned it, and those who have never tried it. These distinctions are important for informing development and marketing strategies (Rogers et al., 2014). Moreover, these studies often focus on usage patterns or user perceptions without connecting them to concrete feature development decisions. As a result, we know little about how specific user needs translate into buildable and valuable AI features.
@@
To address existing knowledge gaps, we conducted a multi-method study in two phases: qualitative semi-structured interview (Kallio et al., 2016) and prediction market (Wolfers and Zitzewitz, 2004). First, we interviewed 35 tech creators across three usage groups (Adopters, Churners, and Non-users) to gain qualitative insights into their experiences, challenges, and unmet needs with current AI features. This phase reveals unmet needs and expectations across five areas: Technology Improvement, Interaction, Customization, Simplifying Skill Building, and Programming Tasks. Yet, understanding user needs alone does not directly translate into feasible product development strategies. Prioritizing feature development requires additional insight into both technical feasibility and organizational readiness. To achieve this, we complemented interviews with a prediction market. Participants predicted which features would be implemented, using collective judgment to estimate outcomes. We asked 102 internal practitioners to predict the likelihood of implementing 12 user-desired features. This allows us to integrate user-driven insights with practical considerations from the very people who shape our product development. Our results reveal that internal practitioners correctly anticipate the implementation of features around AI in Implementation and Context-aware Technology. In contrast, features requiring proactive behavior and maintenance support remain undervalued and underdelivered. These findings highlight a misalignment between developer needs and current feasibility assessments, pointing to concrete opportunities for more helpful and forward-looking AI assistance in IDEs.
@@
With this work, we aim to provide industry practitioners and researchers with structured insights to guide product decisions. Specifically, offer:
@@
  • A categorization of user needs for in-IDE HAX to outline requirements to make AI tools more useful;
@@ 2. BACKGROUND
AI-powered programming assistants have become increasingly common, with tools such as GitHub Copilot, Cursor, and JetBrains AI integrated into developer workflows (12; 16; Cursor AI (2024)). While their technical capabilities continue to evolve, research into how developers experience and evaluate such tools has also grown.
Recent studies have investigated what professional developers need and expect from AI-powered programming assistants (Liang et al., 2024; Mcnutt et al., 2023; Liu et al., 2024; Sergeyuk et al., 2024; Vaithilingam et al., 2022; Zhang et al., 2023; Wang et al., 2023). Across contexts, developers consistently seek tools that reduce repetitive effort and provide relevant suggestions without introducing new debugging burdens (Liang et al., 2024; Zhang et al., 2023; Vaithilingam et al., 2022). Effective assistance is expected to align with the current task, understand code context, and support natural language interaction (Liang et al., 2024; Liu et al., 2024; van Dam et al., 2023). Transparency and control are also emphasized, including the ability to steer and correct outputs and control context (Mcnutt et al., 2023; Liang et al., 2024). Several studies highlight that experienced users tend to expect more than raw generation (Liu et al., 2024; Mcnutt et al., 2023). For example, tools are valued when they adapt to user intent and surface relevant, task-specific information (Liang et al., 2024; Liu et al., 2024; de Moor et al., 2024). At the same time, integration challenges persist. Developers often report friction when tools disrupt their workflow or fail to fit specific environments (Sergeyuk et al., 2024; Vaithilingam et al., 2022). Moreover, non-functional requirements such as performance and scalability may be harder to maintain (Sergeyuk et al., 2024). These insights point to the need for assistants that can flexibly support diverse roles, offer customization, and operate within existing development processes (Mcnutt et al., 2023; Sergeyuk et al., 2024; Liang et al., 2024).
@@
However, most studies examine developer interaction with AI in broad settings, without addressing the unique characteristics of IDE-based workflows. This distinction is important. According to large-scale industry reports, up to 97% of developers rely on IDEs for software tasks (28). In-IDE AI assistants operate within a structured environment that offers richer contextual signals such as file scope, project layout, and cursor history. These tools have the potential to deliver more context-aware support while minimizing context switching. Yet, their design introduces new constraints, including the need to remain non-intrusive and responsive in real time. Therefore, in-IDE AI tools shape distinct user needs, expectations, and challenges that warrant focused investigation.
@@
Moreover, existing studies tend to treat developers as a single user group, overlooking differences across adoption stages (Sergeyuk et al., 2025). However, industry practice suggests that design and feature prioritization must account for the heterogeneous needs of distinct groups: Adopters, Churners, and Non-Users (Rogers et al., 2014; Amershi et al., 2019). These groups not only engage (or not) with AI tools differently, but also hold divergent mental models, trust levels, and expectations, which may lead to different product requirements or marketing strategies. Recognizing these “personas” is necessary to design AI assistants that support retention, re-engagement, or onboarding.
@@
At the same time, even when user needs are clearly articulated, translating them into product decisions remains a challenge. Feature roadmaps must also reflect internal constraints such as technical feasibility, resource allocation, and infrastructure readiness. Traditional prioritization techniques, such as expert panels and surveys, often fall short by providing limited collective insights or biased individual opinions (Avella, 2016). An alternative approach, prediction markets, has been shown to effectively aggregate collective expert judgment to forecast the likelihood of specific outcomes (Wolfers and Zitzewitz, 2004).
A prediction market is a structured method wherein participants place bets on future events, effectively leveraging the “wisdom of the crowd” to estimate probabilities accurately (Wolfers and Zitzewitz, 2004; Arrow et al., 2008). Previously, prediction markets have been successfully used to forecast a broad range of outcomes from presidential votes (Berg and Rietz, 2014), to replicability of scientific research (Dreber et al., 2015), and the future of real products on the market (Bryant and Jensen, 1994). Google has been a pioneer in using internal prediction markets for software engineering and product development decisions. Research by Cowgill and Zitzewitz (2015) analyzed internal prediction markets of several companies including Google, finding they produced more accurate forecasts than experts about company-specific events like product launch dates and business metrics. While prediction markets have shown promise in corporate technology settings, academic research on their application to software engineering research and practice remains limited. This gap represents a significant opportunity for HCI and software engineering research to examine how prediction markets can be adapted for the specific technical and social contexts of modern software development.
@@
Thus, this paper combines qualitative interviews and prediction market, integrating rich user insights with realistic assessments from internal stakeholders. This dual perspective enables a grounded understanding of both user needs and internal feasibility, offering a practical contribution to the design and deployment of in-IDE AI assistants. We address the following research questions:
  • RQ1. What are developers’ needs when using in-IDE AI tools?
@@
  • RQ3. To what extent do practitioners acknowledge and correctly anticipate which AI features inspired by user needs are likely to be implemented?
3. UNDERSTANDING DEVELOPER NEEDS
In this section, we examine how developers experience AI inside their IDEs when coding. Our goal is to provide a structured understanding of what developers need from AI assistance, what challenges they face, and how these vary across adoption groups. We address RQ1 and RQ2 through qualitative interviews.
3.1. METHOD
@@
The study was conducted in May-June 2024 in line with our institutional ethical standards, adhering to the values and guidelines outlined in the ICC/ESOMAR International Code (ICC/ESOMAR, 2024).
3.1.1. Recruitment
@@
We recruited participants from users of JetBrains, a software engineering company that also develops AI-based tools for coding, to ensure a consistent experience with development environments. Note that with that, participants could use any in-IDE AI assistant. Invitations were sent via email to consenting subscribers and promoted through the authors’ social media.
Respondents completed a qualification survey. It was designed to assign them to one of three groups based on a question about AI tool experience: current users were classified as Adopters, former users as Churners, and those who never used such tools as Non-users. We chose these three groups to capture various experiences with AI tools in IDEs (Rogers et al., 2014). These groups are referred to using the prefixes A, C, and N, respectively, in participant identifiers throughout the paper. In the survey, we collected data on job roles, experience levels, IDEs used, AI tools tried for coding, and duration of AI usage. Out of 380 responses, we iteratively invited participants until data saturation was reached (Hennink and Kaiser, 2022), which means that no new themes or insights emerged from the last interviews. As a result, we conducted 35 interviews: 15 Adopters, 12 Churners, and 8 Non-Users. Most of the participants were developers (32) with overlapping roles: 9 team leads, 8 architects, and 5 DevOps engineers. The sample included 28 seniors, 5 mid-level practitioners, and 2 juniors, with experience ranging from under 5 years (9) to over 11 years (16). Participants chose a reward between a USD 100 Amazon gift card or an equivalent-value pack of all JetBrains products.
3.1.2. Data Collection
@@
The interviews followed a semi-structured script developed by the first two authors, reviewed by certified UX researchers at JetBrains, and piloted with four internal participants (Kallio et al., 2016). The script included open-ended questions on AI experience, utilized tools, challenges and successes with in-IDE HAX, unmet needs, desired features, and visions for AI in coding.
@@
All interviews were conducted remotely by one of the first two authors, lasting 71 minutes on average (range: 46-92). Each interview followed a common structure covering challenges, successes, needs, and future expectations, with question framing adapted to each group’s AI experience while maintaining cross-group comparability. For Adopters, the questions focused on their ongoing use of AI tools, exploring typical workflows, challenges, and successes. Churners, those who have previously used AI tools but then stopped using them, reflected on their initial adoption and further disengagement, identifying reasons for it and missed opportunities. Non-Users were asked about their reasons for not adopting AI, any potential barriers and possible motivations that could encourage them to begin using these tools. This enabled group-specific depth while maintaining cross-group comparability during analysis.
3.1.3. Data Analysis
@@
All interviews were transcribed using Condens [^1] Condens: a user research repository for analyzing qualitative data. https://condens.io/. We applied a qualitative content analysis (Forman and Damschroder, 2007; Elo and Kyngäs, 2008) to identify user-reported successes, challenges, unmet needs, and future expectations for in-IDE AI. The first two authors reviewed the first transcript together to agree on the analysis and the code-deriving process. They then went into an iterative inductive process, where they independently assigned Codes to quotes from the transcripts. Eventually, they compiled a list of 161 codes, met again to add or remove codes, and resolved any disagreements until full consensus was reached. The final versions of the codes were then assigned to the corresponding quotes. The codes were then grouped into Thematic groups representing key areas of need. These were organized into broader Topics that describe the categories of in-IDE HAX needs. For example, the following quote:
@@
was coded as “Code Review Support”, assigned to the group “Code Optimization”, under the topic “Simplifying Programming Tasks”. Following established practices, we identified emergent patterns across user groups, recognizing that in qualitative research, the presence and salience of themes matter more than frequency counts.
@@
All materials used for data gathering, together with a full demography table of participants and a summarized codebook, are included in Supplementary Materials (JetBrains Research, 2025). Full interview transcripts cannot be shared due to the company’s confidentiality policy.
3.2. FINDINGS
@@
Figure 1. The Categorization of Users’ Needs in HAX Within IDEs With Descriptions
. A hierarchical taxonomy diagram presenting 25 user needs for AI assistants in IDEs, organized into five color-coded categories. Each need includes a title, description, and example features. Technology Improvements (orange header) contains nine needs: Accessible Technology (craft affordable and adaptable technologies; examples: separate feature purchase), Proactive Technology (reduce manual effort and preemptively solve issues; examples: proactive optimization, AI performing actions), Reliable Technology (ensure robustness and accuracy to increase trust; examples: output clarifications, providing sources), Up-to-date Technology (integrate latest advancements; examples: SOTA tooling, up-to-date data), Ethical Technology (balance innovation with conscientious practices; examples: rules and restrictions, automated regulation), Private Technology (safeguard privacy and data security; examples: on-premise, local AI), Secure Technology (fortify against vulnerabilities; examples: annotating AI code), Aligned Technology (ensure coherent, non-redundant solutions consistent with codebase; examples: library suggestions, better commit messages), and Context-aware Technology (understand and integrate various context levels; examples: manually set context, project-wide awareness). Simplifying Programming Tasks (purple header) contains five needs: AI Support in SDLC (universal adoption for tedious tasks across development lifecycle; examples: universal programming assistance), Navigating the Codebase (facilitate deeper code comprehension; examples: onboarding support, database analytics), AI in Ideation (enhance creative brainstorming; examples: creative AI, cross-model prompting), AI in Implementation (support development workflows from concept to implementation; examples: multi-file generation, project setup assistance), AI in Testing (enhance test generation and evaluation; examples: coverage report, UI testing), AI in Maintenance (simplify optimization and ongoing updates; examples: repository and dependency management), Code Optimization (enhance code efficiency and quality; examples: code evaluation, code review support), and External System Integration (incorporate AI into broader workflows and CI/CD pipelines; examples: AI in CI/CD pipelines, out-of-IDE AI access). Technology Interaction (pink header) contains two needs: Non-Interruptive Technology (maintain smooth workflow by reducing interruptions; examples: stable AI functionality, voice interaction) and Internal System Integration (enhance cohesion of AI within development environments; examples: anywhere-in-IDE AI access). Technology Customization (teal header) contains two needs: Adjustable Technology (configurable AI adapting to diverse requirements; examples: configurable AI, AI action customization) and Optimizing Workflow (create cohesive work environment supporting daily tasks; examples: daily ecosystem summarizer). Simplifying Skill Building (green header) contains two needs: User Education (equip users with knowledge to effectively interact with AI; examples: shareable prompt library) and Learning Support (support acquisition of new concepts and skills; examples: step-by-step output revealing).
3.2.1. Categorization of Developer Needs